# FIPO: Free-form Instruction-oriented Prompt Optimization with Preference Dataset and Modular Fine-tuning Schema

### v3

@@ Abstract
When the quality of naive prompts is carefully optimized by human experts, the task performance of large language models (LLMs) can be significantly improved. However, expert-based prompt optimizations are expensive. Herein, some works have proposed Automatic Prompt Optimization (APO), to optimize naive prompts according to task outputs of given in-box testing models, with the help of advanced LLMs (e.g., GPT-4) in an ad-hoc way. Although effective, existing schemes suffer from poor generalization ability and privacy risk. To this end, we collect the first large-scale Prompt Optimization Preference dataset (POP), fine-tune offline local LLM-based optimizers, then fairly test with various downstream models. Our method allows accurate optimization of the core task instruction part within the naive prompt in a model-agnostic manner, and thus is named Free-from Instruction-oriented Prompt Optimization (FIPO). In specific, FIPO uses a modular APO template that dynamically integrate the naive task instruction, optional instruction responses, and optional ground truth to produce finely optimized prompts. The POP dataset is meticulously constructed using advanced LLMs, undergoing rigorous cross-validation by human experts and analytical models. Leveraging insights from the data with Tulu2 models and diverse fine-tuning strategies, we validate the efficacy of FIPO framework across five public benchmarks and six testing models [^1] Check codes and data here: https://github.com/LuJunru/FIPO_Project..
@@
Junru Lu1, Siyu An2, Min Zhang3, Yulan He4, Di Yin2, Xing Sun2 1University of Warwick, 2Tecent YouTu Lab, 3Westlake University, 4King’s College London 1junru.lu@warwick.ac.uk, 3zhangmin@westlake.edu.cn, 4yulan.he@kcl.ac.uk 2{siyuan, endymecyyin, winfredsun}@tencent.com
@@ 1 Introduction
Large Language Models (LLMs) show their intelligence Zhao et al. (2023a); Yang et al. (2023c); Achiam et al. (2023) on various benchmarks Cobbe et al. (2021); Suzgun et al. (2023); Bisk et al. (2020); Huang et al. (2019); Hendrycks et al. (2021). However, the task performance of LLMs depends on the given task prompt. LLMs may not be able to give the correct answer based on the naive task prompt, but after careful optimization by human experts, it can perform the same task well under the guidance of high-quality prompts Wei et al. (2022); Kojima et al. (2022); Yang et al. (2023b).
@@
Figure 1: Online ad-hoc APO vs. our Local End2end FIPO. Although both leverage advanced LLMs (e.g., GPT-4), in FIPO, we design a local training pipeline and completely removed the reliance on any in-box model generators.
@@
Obviously, expert-based prompt optimizations are expensive. Therefore, in recent years, Automatic Prompt Optimization (APO) has emerged as a prominent area of research. Discrete APO is one of the popular strategies, focusing on identifying optimal combinations of discrete tokens to serve as optimized prompts van de Kar et al. (2022); Yuan et al. (2021); Jiang et al. (2020); Pryzant et al. (2023). Particularly, there has been significant interest in LLM-based discrete APO Zhou et al. (2023); Do et al. (2023); Wang et al. (2023a), which introduce ad-hoc strategies leveraging leading API-accessible LLMs (e.g., GPT-4 Achiam et al. (2023)).
@@
These APO approaches typically involve iterative optimization between an in-box testing generator $\bm{M}_{g-in}$ and an advanced optimizer $\bm{M}_{o-api}$. As illustrated in the upper half of Figure 1, the generator $\bm{M}_{g-in}$ first provide answers upon the naive prompt $\bm{x}^{n}$: “Calculate the average value of the list”, and then the optimizer $\bm{M}_{o-api}$ provides rational feedback and several upgraded prompt candidates $\{\bm{x}^{o}\}$. This above process can be repeated several times. Afterwards, a high-quality optimized prompt $\bm{x}^{o}$ is finally tailored for the in-box generator $\bm{M}_{g-in}$, which can be ensured to produce the desired response: “The answer value is 44.25”. However, despite effective, several drawbacks exist: (1) Privacy risk. The entire online optimization is exposed to external LLM services, which poses a privacy risk; (2) Poor Generalization Ability. The ad-hoc optimization is model-centric, as it runs with instant testing responses from the in-box generator $\bm{M}_{g-in}$, leading to performance degradation when tested with out-of-box generators $\bm{M}_{g-out}$. For instance, a wrong answer is inferred by the out-of-box generator $\bm{M}_{g-out}$: “44.1”.
@@
To address the existing challenges, we introduce Free-form Instruction-oriented Prompt Optimization (FIPO). The bottom half of Figure 1 illustrates our FIPO. Unlike the online ad-hoc APO approach, FIPO directly fine-tunes the general local optimizer $\bm{M}_{local}$, and apply over any testing generator $\bm{M}_{g}$. In specific, we first design a meta template towards universal APO (Figure 2), and thereby collect 30,000 prompt optimization preference examples from advanced optimizer $\bm{M}_{o-api}$ (§ 2.3). We demonstrate the reliability of this data through multiple cross-validation methods (Table 1). On this basis, we then apply with mainstream end-to-end fine-tuning strategies for obtaining a valid local optimizer $\bm{M}_{o-local}$ (Figure 3). Particularly, in the Figure 1’s example, our local optimizer $\bm{M}_{local}$ provides insightful hints in the optimized prompt $\bm{x}^{o}$: “calculate the average value by finding the sum of all elements and dividing it by the total number of elements in the list”. Such stepwise guidance generally enhance the answer quality of any testing generator $\bm{M}_{g}$. Our contributions are as follows:
@@
    We denote the drawbacks of previous online ad-hoc APO, and then introduce our local free-form instruction-oriented prompt optimization (FIPO). We design a generic APO template that allows the optimizer to unpack the pre-built testing generators, which is a key component for former methods.
@@
    Based on the template, we compile a 30k large-scale Prompt Optimization Preference dataset (POP), and explore with several mainstream fine-tuning strategies: Supervised Fine-tuning (SFT), Direct Preference Optimization (DPO) Rafailov et al. (2023) and Identity Preference Optimization (IPO) Azar et al. (2023). We further develop an Iterative Preference Learning (IPL) pipeline for self-enhanced preference optimization.
@@
    We evaluate the efficacy and adaptability of fine-tuned FIPO optimizer across five downstream benchmarks and three diverse generators, and obtain superior results compared to previous online ad-hoc APO methods.
Figure 2: Step 1 and 2 of FIPO: (1) Design a meta template towards universal APO; (2) Collect 30,000 large-scale prompt optimization preference exemplars from a suboptimal LLM GPT-3.5-turbo and an optimal LLM GPT-4.
@@ 2 Methodology
In this section, we start with task formulation (§ 2.1), then introduce our meta template towards universal APO (§ 2.2), collected POP data (§ 2.3) and training strategies (§ 2.4).
@@ 2.4 Fine-tuning Strategies
Dataset Diversification. In the left hand of Figure 3, we evenly split the 30k samples as eight types depending on the existence of naive response $\bm{\hat{y}}^{n}$ and ground truth response $\bm{y}^{n}$, as well as a format condition “generation” or “multi-choice”. Directionally fine-tuning optimizer has to relay on pre-generated responses, while no any response will be exposed during inference. Hence, the dataset diversification is necessary to help reduce the exposure gap between training and testing, and generalize the original “generation” instruction format to another common “multi-choice” instruction format. The left bottom corner in Figure 3 takes an example of Type 6. The responses $\bm{\hat{y}}^{n}$ and $\bm{y}^{n}$ are modified as candidates adhere to the naive prompt $\bm{x}^{n}$. We then set “A” and “B” as new naive response $\bm{\hat{y}}^{n}$ and ground truth response $\bm{y}^{n}$ [^3] More details of data diversification are in Appendix B..
@@
where $G(*)$ denotes as either IPO or DPO loss [^4] Algorithmic details of IPL lie in Appendix B and  C..
@@ 3 Experiments
The ultimate target of FIPO lies in the general performance enhancement with downstream generators $\bm{M}_{g}$ (Eq. 3). Herein, in the evaluation, we first use fine-tuned optimizer $\bm{M}_{o}$ to produce optimized testing prompts $\bm{\hat{x}}_{t}^{o}$, then obtain optimized testing response $\bm{\hat{y}}_{t}^{o}$ and naive testing response $\bm{\hat{y}}^{n}_{t}$ for answer quality checking, as shown in the bottom right corner of Figure 3. We begin with our experimental settings (§ 3.1), efficacy presentation and comparisons against online ad-hoc APO methods (§ 3.2.1), then follow with analysis of different fine-tuning strategies (§ 3.2.2), and case analysis (§ 3.2.3) [^5] Hyperparameters and training cost are in Appendix A..
@@ 3.1.2 Foundation models
Figure 4: The FIPO-optimized prompts help various downstream testing LLMs (X-axis) gain more promising improvements, compared with other prompt optimization approaches (shown by the bars). We specifically annotate the improvements of FIPO against naive prompts from the original dataset (). More details: Appendix G and  D.
@@ 3.2.2 Fine-tuning Analysis
Dataset diversification is necessary. In the bottom of Table 3, we present ablation studies of preprocessing dataset diversification, mentioned in section 2.4. In specific, IPO-70B-gen stands for not diversifying half of the training set into multi-choice format, which is introduced as type 5,6,7 and 8 in Figure 3. As for IPO-70B-partial, we only use type 3,4,7 and 8 in Figure 3 as pairwise diversification templates. The ablated optimizer weakens all benchmarks, except BBH, which is due to its unique symbolic reasoning pattern (Appendix E).
@@
In Table 4, we present several examples from the downstream testing benchmarks, discussing the efficacy and shortcomings of FIPO. Particularly, we smear (key optimized content with blue), and overwhelmed (cheating notes with underlines). The 1st optimized prompt of BBH case explicitly mentions that ”2000 is a leap year”, which is a key detail for calculating dates in February. The 2nd optimized prompt of MMLU question capitalizes “NOT” to draw attention to the negative aspect, ensuring the model focuses on identifying the incorrect option. It also explicitly instructs the model to provide the letter of the incorrect option, reducing ambiguity. And the 3rd optimized prompt of CosmosQA case provides definitions for “miss” and “hit” according to Signal Detection Theory, making it easier for the model to understand and apply the correct term. While in the last GSM8K case, FIPO breaks down the calculation into clear, step-by-step instructions, ensuring the model understands the process of finding the average. However, it provides overwhelmed cheating notes of the final answer. We manually confirm that this issue mostly occurs in GSM8K and BBH’s math calculation questions, happening <10%, in Appendix H.
@@ 4 Related Work
Preference Optimization for LLMs allows LLMs to align with human minds in a more nuanced way Pan et al. (2023b), compared to SFT. Proximal Policy Optimization (PPO) is one of the well-known preference optimization approach to first train a reward model with pairwise human preference data, then align LLMs with the reward model via reinforcement learning Ouyang et al. (2022); Bai et al. (2022). Despite its efficacy, PPO is often blamed for its training instability and expensive costs. To this end, Direct Preference Optimization (DPO) have been proposed, aiming to align LLMs through implicit modeling, thereby eliminating the flaws associated with the explicit use of reward models Rafailov et al. (2023). A series of following work of DPO are further presented Azar et al. (2023); Zhao et al. (2023b); Ethayarajh et al. (2024).
@@ 5 Conclusion
  • Overwhelmed Cheating Notes. As observed in the case analysis, FIPO sometimes provides overly detailed instructions that can be considered as “cheating notes”. This issue is particularly prevalent in tasks involving mathematical calculations. While this enhances performance, it may not align with the intended use of prompt optimization.
  • Evaluation Metrics. The current evaluation primarily focuses on accuracy metrics. While accuracy is important, other aspects such as the interpretability, fairness, and ethical implications of the optimized prompts should also be considered in future work.
  • Optimization of In-context Examplars. Our current research on FIPO does not include optimization of the in-context examples, but rather focuses on prompt optimization for the task instructions.
Appendix A HyperParameters and Training Cost
We report hyperparameters and training cost in Table 5. We fine-tune all models on same computational node server equipped with 900G CPU RAM and 8 X A100 40G GPUs. We only use single node for fine-tuning 13B models, but use 4 nodes for 70B models. While the batch size varies, we maintain a consistent global batch size on every node. In terms of different fine-tuning strategies, most hyperparameters are shared. The only difference lies on the learning rate, as we find large learning rate for DPO and IPO will cause collapse. What’s more, all preference learning approaches relay on one external hyperparameter $\bm{\beta}$. We set its value as 0.01 empirically Tunstall et al. (2023).

### v4

@@ Abstract
When carefully optimized by human experts, naive prompts can significantly enhance the task performance of large language models (LLMs). However, such expert-driven prompt optimizations are resource-intensive. To address this, some studies have proposed Automatic Prompt Optimization (APO), which refines naive prompts according to task outputs from in-box testing models, utilizing advanced LLMs (e.g., GPT-4) in an ad-hoc way. Although effective, current approaches face challenges in generalization and privacy risks. To overcome these limitations, we have developed the first large-scale Prompt Optimization Preference (POP) dataset, fine-tuned offline local LLM-based optimizers, and conducted fairly evaluations across various downstream models. Our method, named Free-from Instruction-oriented Prompt Optimization (FIPO), allows precise optimization of the core task instructions in naive prompts in a model-agnostic manner. FIPO uses a modular APO template that dynamically incorporates the naive task instructions, optional instruction responses, and optional ground truth to produce refined prompts. The POP dataset is meticulously constructed using advanced LLMs, undergoing rigorous cross-validation by human experts and analytical models. By leveraging insights from this dataset, along with Tulu2 models and diverse fine-tuning strategies, we validate the efficacy of the FIPO framework across five public benchmarks and six testing models. [^1] Our dataset and codes are available at: https://github.com/LuJunru/FIPO_Project.
@@
Junru Lu1,2, Siyu An2, Min Zhang3, Yulan He1,4,5, Di Yin2, Xing Sun2 1University of Warwick, 2Tecent YouTu Lab, 3East China Normal University, 4King’s College London, 5The Alan Turing Institute 1junru.lu@warwick.ac.uk, 3mzhang@cs.ecnu.edu.cn, 4yulan.he@kcl.ac.uk 2{siyuan, endymecyyin, winfredsun}@tencent.com
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities Zhao et al. (2023a); Yang et al. (2023c); Achiam et al. (2023) across various benchmarks Cobbe et al. (2021); Suzgun et al. (2023); Bisk et al. (2020); Huang et al. (2019); Hendrycks et al. (2021). However, their task performance is highly dependent on the quality of the given task prompt. While LLMs may struggle to produce correct answers when working with naive task prompts, they can excel on the same tasks when guided by carefully optimized, high-quality prompts crafted by human experts Wei et al. (2022); Kojima et al. (2022); Yang et al. (2023b).
@@
Figure 1: Online Ad-hoc APO vs. our Local End-to-End FIPO: Although both approaches leverage advanced LLMs (e.g., GPT-4), FIPO introduces a a locally trained pipeline that eliminates any dependence on in-box model generators, ensuring a fully self-contained and end-to-end optimization process.
@@
Obviously, expert-based prompt optimization is costly. Therefore, in recent years, Automatic Prompt Optimization (APO) has emerged as a prominent area of research. Discrete APO is one of the popular strategies, focusing on identifying optimal combinations of discrete tokens to serve as optimized prompts van de Kar et al. (2022); Yuan et al. (2021); Jiang et al. (2020); Pryzant et al. (2023). Particularly, there has been significant interest in LLM-based discrete APO Zhou et al. (2023); Do et al. (2023); Wang et al. (2023a), which introduce ad-hoc strategies leveraging leading API-accessible LLMs (e.g., GPT-4 Achiam et al. (2023)).
@@
These APO approaches typically involve iterative optimization between an in-box testing generator $\bm{M}_{g-in}$ and an advanced optimizer $\bm{M}_{o-api}$. As illustrated in the upper half of Figure 1, the generator $\bm{M}_{g-in}$ first responds to a naive prompt $\bm{x}^{n}$ such as: “Calculate the average value of the list”, and then the optimizer $\bm{M}_{o-api}$ provides rational feedback and suggests several upgraded prompt candidates $\{\bm{x}^{o}\}$. This iterative process continues until a high-quality optimized prompt $\bm{x}^{o}$ is generated. This final prompt is tailored to the in-box generator $\bm{M}_{g-in}$, ensuring it produces the desired response, e.g., “The answer value is 44.25”.
Despite their effectiveness, several drawbacks remain: (1) Privacy Risk. The entire online optimization process relies on external LLM services, exposing sensitive information to third-party systems; (2) Poor Generalization. Ad-hoc optimization is highly model-specific, as it depends on immediate testing responses from the in-box generator $\bm{M}_{g-in}$, leading to performance degradation when tested with out-of-box generators $\bm{M}_{g-out}$. For instance, the out-of-box generator $\bm{M}_{g-out}$ might produce an incorrect response like “44.1”.
@@
To address the above limitations, we introduce Free-form Instruction-oriented Prompt Optimization (FIPO). The bottom half of Figure 1 illustrates the FIPO framework. Unlike the online ad-hoc APO approach, FIPO directly fine-tunes a general local optimizer $\bm{M}_{local}$, and applies it across any testing generator $\bm{M}_{g}$. In specific, we first design a meta-template for universal APO (Figure 2), enabling the collection of 30,000 prompt optimization preference examples using an advanced optimizer $\bm{M}_{o-api}$ (§ 2.3). We demonstrate the reliability of this dataset through multiple cross-validation methods (Table 1). Building on this foundation, we employ mainstream end-to-end fine-tuning strategies to create an effective local optimizer $\bm{M}_{o-local}$ (Figure 3). For example, in Figure 1, the local optimizer $\bm{M}_{local}$ generates an optimized prompt $\bm{x}^{o}$, providing clear, stepwise guidance: “calculate the average value by finding the sum of all elements and dividing it by the total number of elements in the list”. Such stepwise guidance significantly enhances the answer quality across any testing generator $\bm{M}_{g}$. Our contributions are as follows:
@@
    We highlight the drawbacks of previous online ad-hoc APO, and introduce FIPO, a local free-form instruction-oriented prompt optimization. FIFO employs a generic APO template that allows the optimizer to operate independently of pre-built testing generators, a limitation of former methods.
@@
    Leveraging the APO template, we compile a large-scale Prompt Optimization Preference dataset (POP) with 30,000 examples. We explore several mainstream fine-tuning strategies, including Supervised Fine-tuning (SFT), Direct Preference Optimization (DPO) Rafailov et al. (2023), and Identity Preference Optimization (IPO) Azar et al. (2023), while also developing an Iterative Preference Learning (IPL) pipeline for enhanced preference optimization.
@@
    We evaluate the efficacy and adaptability of the fine-tuned FIPO optimizer across five downstream benchmarks and three diverse generators, and obtain superior results compared to existing online ad-hoc APO methods.
Figure 2: Step 1 and 2 of FIPO: (1) Design a meta-template for universal APO; (2) Collect 30,000 large-scale prompt optimization preference exemples using a suboptimal LLM (GPT-3.5-turbo) and an optimal LLM (GPT-4).
@@ 2 Methodology
In this section, we start with task formulation (§ 2.1), then introduce our meta-template for universal APO (§ 2.2), the collected POP data (§ 2.3) and the training strategies employed (§ 2.4).
@@ 2.4 Fine-tuning Strategies
Dataset Diversification. In the left hand of Figure 3, we evenly split the 30k samples as eight types depending on the existence of naive response $\bm{\hat{y}}^{n}$ and ground truth response $\bm{y}^{n}$, as well as a format condition “generation” or “multi-choice”. Directionally fine-tuning optimizer has to relay on pre-generated responses, while no any response will be exposed during inference. Hence, the dataset diversification is necessary to help reduce the exposure gap between training and testing, and generalize the original “generation” instruction format to another common “multi-choice” instruction format. The left bottom corner in Figure 3 takes an example of Type 6. The responses $\bm{\hat{y}}^{n}$ and $\bm{y}^{n}$ are modified as candidates adhere to the naive prompt $\bm{x}^{n}$. We then set “A” and “B” as new naive response $\bm{\hat{y}}^{n}$ and ground truth response $\bm{y}^{n}$ [^3] More details of data diversification are in Appendix A..
@@
where $G(*)$ denotes as either IPO or DPO loss [^4] Algorithmic details of IPL lie in Appendix A and  C..
@@ 3 Experiments
The ultimate target of FIPO lies in the general performance enhancement with downstream generators $\bm{M}_{g}$ (Eq. 3). Herein, in the evaluation, we first use fine-tuned optimizer $\bm{M}_{o}$ to produce optimized testing prompts $\bm{\hat{x}}_{t}^{o}$, then obtain optimized testing response $\bm{\hat{y}}_{t}^{o}$ and naive testing response $\bm{\hat{y}}^{n}_{t}$ for answer quality checking, as shown in the bottom right corner of Figure 3. We begin with our experimental settings (§ 3.1), efficacy presentation and comparisons against online ad-hoc APO methods (§ 3.2.1), then follow with analysis of different fine-tuning strategies (§ 3.2.2), and case analysis (§ 3.2.3) [^5] Hyperparameters and training cost are in Appendix B..
@@ 3.1.1 Baselines
Figure 4: The FIPO-optimized prompts help various downstream testing LLMs (X-axis) gain more promising improvements, compared with other prompt optimization approaches (shown by the bars). We specifically annotate the improvements of FIPO against naive prompts from the original dataset (). More details: Appendix G and  7.
@@ 3.2.2 Fine-tuning Analysis
Dataset diversification is necessary. In the bottom of Table 3, we present ablation studies of preprocessing dataset diversification, mentioned in section 2.4. In specific, IPO-70B-gen stands for not diversifying half of the training set into multi-choice format, which is introduced as type 5,6,7 and 8 in Figure 3. As for IPO-70B-partial, we only use type 3,4,7 and 8 in Figure 3 as pairwise diversification templates. The ablated optimizer weakens all benchmarks, except BBH, which is due to its unique symbolic reasoning pattern (Appendix F).
@@
In Table 4, we present several examples from the downstream testing benchmarks, discussing the efficacy and shortcomings of FIPO. Particularly, we smear (key optimized content with blue), and overwhelmed (cheating notes with underlines). The 1st optimized prompt of BBH case explicitly mentions that ”2000 is a leap year”, which is a key detail for calculating dates in February. The 2nd optimized prompt of MMLU question capitalizes “NOT” to draw attention to the negative aspect, ensuring the model focuses on identifying the incorrect option. It also explicitly instructs the model to provide the letter of the incorrect option, reducing ambiguity. And the 3rd optimized prompt of CosmosQA case provides definitions for “miss” and “hit” according to Signal Detection Theory, making it easier for the model to understand the correct term. While in the last GSM8K case, FIPO breaks down the calculation into clear, step-by-step instructions, ensuring the model understands the process of finding the average. However, it provides overwhelmed cheating notes of the final answer [^6] This issue mostly occurs in GSM8K and BBH’s math calculation questions, happening ¡10%, in Appendix H..
@@ 4 Related Work
Preference Optimization for LLMs allows LLMs to align with human minds in a more nuanced way Pan et al. (2023b), compared to SFT. Proximal Policy Optimization (PPO) is one of the well-known preference optimization approach to first train a reward model with pairwise human preference data, then align LLMs with the reward model via reinforcement learning Ouyang et al. (2022); Bai et al. (2022). Despite its efficacy, PPO is often blamed for its training instability and expensive costs. To this end, Direct Preference Optimization (DPO) have been proposed, aiming to align LLMs through implicit modeling, thereby eliminating the flaws associated with the explicit use of reward models Rafailov et al. (2023). Following works of DPO are presented Azar et al. (2023); Zhao et al. (2023b); Ethayarajh et al. (2024); Lu et al. (2024).
@@ 5 Conclusion
(1) Overwhelmed Cheating Notes. As shown in the case analysis, FIPO sometimes provides overly detailed instructions that can be considered as “cheating notes”. This issue is particularly prevalent in tasks involving mathematical calculations. While this enhances performance, it may not align with the intended use of prompt optimization. (2) Evaluation Metrics. The current evaluation primarily focuses on accuracy metrics. While accuracy is important, other aspects such as the interpretability, fairness, and ethical implications of the optimized prompts should also be considered in future work. (3) Optimization of In-context Examplars. FIPO does not include optimization of the in-context examples, but rather focuses on the optimization for the task instructions only.
Acknowledgements
This work was supported in part by the UK Engineering and Physical Sciences Research Council (EPSRC) through a Turing AI Fellowship (grant no. EP/V020579/1, EP/V020579/2), and Innovate UK through the Accelerating Trustworthy AI programme (grant no. 10093055).
Appendix A Dataset Diversification
Appendix B HyperParameters and Training Cost
We report hyperparameters and training cost in Table 5. We fine-tune all models on same computational node server. We only use single node for fine-tuning 13B models, but use 4 nodes for 70B models. While the batch size varies, we maintain a consistent global batch size on every node. In terms of different fine-tuning strategies, most hyperparameters are shared. The only difference lies on the learning rate, as we find large learning rate for DPO and IPO will cause collapse. What’s more, all preference learning approaches relay on one external hyperparameter $\bm{\beta}$. We set its value as 0.01 empirically Tunstall et al. (2023).